OpenStudy (anonymous):

PLEASE HELPPPPPPP ME!!!!!! I WILL THANK YOU THIS FOREVR :D Sasha finished mowing lawns at 4:15 PM. It took 1 3/4h to mow the first lawn. The second and third lawns she mowed each took 1 1/5h. She took only a 45-min break. When did Sasha begin mowing the first lawn? (2 points)10:05 A.M. 11:05 A.M. 11:21 A.M. 12:21 P.M.

OpenStudy (anonymous):

@mark_o. @mathstudent55 HELPPPPP MEEEEE!

OpenStudy (helder_edwin):

1 h = 60 min

OpenStudy (helder_edwin):

1/5 h = 1/5 h 60 min/h=60/5 min = 12 min

OpenStudy (helder_edwin):

so 1 1/5 h = 1 h + 12 min

OpenStudy (helder_edwin):

do the same with the remaining data

OpenStudy (helder_edwin):

did u understand?

OpenStudy (anonymous):

kINDA

OpenStudy (helder_edwin):

3/4 h =

OpenStudy (helder_edwin):

finish this

OpenStudy (helder_edwin):

3/4 h = 3/4 h 60 min/h=

OpenStudy (anonymous):

I DONT KNOW IM NOT SMART! PLEASE HELP ME!!

OpenStudy (nubeer):

\[\frac{ 3}{ 4 } \times60 \min \]

OpenStudy (nubeer):

thts what he is saying

OpenStudy (helder_edwin):

3/4 h = 60(3/4) min = 45 min

OpenStudy (anonymous):

SO WHAT THE ANSWER?

OpenStudy (helder_edwin):

now add up 1 3/4 h = 1 h + 45 min first 1 1/5 h = 1 h + 12 min second 1 1/5 h = 1 h + 12 min third 45 min = 45 min break ________________________ 3 h + 114 min total

OpenStudy (helder_edwin):

but 1 h = 60 min so 3 h + 114 min = 4 h + 54

OpenStudy (anonymous):

IM SO CONFUSED CAN YOU JUST TELL ME WHAT THE ANSWER IS A B C OR D PLEASE

OpenStudy (helder_edwin):

4 h 15 pm = 16 h 15 minus 4 h 54 ----------------------- 11h 21 this the starting hour

OpenStudy (anonymous):

oMG IM REALLY SORRYYY THANK YOU SO MUCHOOOOOOO!!! I THOUGHT YOU WERE GOING TO GET MAD AT ME SORRY AND THANK YOU

OpenStudy (helder_edwin):

u r welcome